40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 24 19:42:12 2014, Charles, Frogs, VAC, Power Failure 
    Reply  Mon Mar 24 21:36:28 2014, Koji, Update, General, Power Failure 
       Reply  Mon Mar 24 22:13:37 2014, steve, Update, General, Power Failure poweroutage.png
          Reply  Tue Mar 25 14:52:57 2014, steve, Update, VAC, vacuum is recovered 17hrsNoPumping.png
             Reply  Fri Mar 28 16:26:20 2014, steve, Update, VAC, RGA scan after power failure afterPowerFailure.png
          Reply  Tue Mar 25 16:11:24 2014, Koji, Update, General, Power Failure 
             Reply  Wed Mar 26 11:30:07 2014, Koji, Update, General, Power Failure 
                Reply  Wed Mar 26 22:22:43 2014, Manasa, Update, General, Recovery 
                   Reply  Thu Mar 27 10:34:39 2014, ericq, Update, General, Recovery 
Message ID: 9755     Entry time: Wed Mar 26 22:22:43 2014     In reply to: 9752     Reply to this: 9756
Author: Manasa 
Type: Update 
Category: General 
Subject: Recovery 

The following that went unnoticed from yesterday were recovered today:

1. ETMX and ETMY 'misalign' scripts weren't running. Troubleshooting showed slow machines c1auxex and c1auxey weren't responding. The machines were reset.

2. PRM oplev gains were zero. Gain values were set looking back at the burt files.

3. X end PZT power supplies were turned ON and set to 100V.

4. X end doubler temperature was reset to the last optimal value on elog (36.35 deg).

 

Some hitches that should be looked into:

1. Check: ASS for X arm seems not quite doing its job. ETMX has to be moved using sliders to obtain maximum TRX and the arm alignment was seen to be drifting.

2. Check: Status of other slow machines and burt restore whichever needs one.

ELOG V3.1.3-